# Qualitatively altered driven Dicke superradiance in extended systems due to infinitesimal perturbations

### v1

@@ IV.2 An angular momentum coupling scheme
For the dependence of the degeneracy on the transmon number, we use the eigenstates of the operators $\{\hat{S}_{1}^{2},\hat{S}_{12}^{2},\hat{S}_{13}^{2},...,\hat{S}_{1N}^{2},\hat{S}_{z}\}$, where $\hat{S}_{1n}^{2}=(\hat{S}_{1}+\hat{S}_{2}+...+\hat{S}_{n})^{2}$. Given $S_{n}=\frac{1}{2}$ and $\hat{S}_{1N}=\hat{S}$, the eigenstates can be written as:
@@
If $a$ or $b$ can take $D_{S}$ values, or the degeneracy for the state |$\ket{S,m}$ is $D_{S}$, there should be $D_{S}^{2}$ possible steady-state density operators associated with the total spin $S$. So the total degeneracy (or the possible number of steady states) $N_{ss}$ is:
$N_{ss}=\sum_{S}D_{S}^{2}$ (35)
By [1, 63, 94], the expression of $D_{S}$ for $N$ transmons with total spin $S$ reads:
@@
Using Eq. (35) and (36), we calculate the degeneracy and compare it to the results from the diagonalization of the Liouvillian in Fig. 10.
Figure 10: Comparison between the degeneracy from the diagonalization of the Liouvillian and the theory from the angular momentum scheme
@@ Appendix A Degeneracy of the Liouvillian’s eigenvalues, counting the number of possible steady states
Because the equation of motion for the perfect driven Dicke model in Eq. (37) and (40) does not depend on the angular momentum coupling, any choice of $a$ and $b$ will share the same coefficients $\rho^{S}_{M,M^{\prime}}$, which is the origin of the degeneracy of the possible steady-state density operators. For a given total spin $S$ and transmon number $N$, the degeneracy of the angular momentum coupling degrees of freedom is[1, 63, 94]:
$D_{S}=(2S+1)\frac{N!}{(\frac{N}{2}+S+1)!(\frac{N}{2}-S)!}$ (43)

### v2

@@ IV.2 An angular momentum coupling scheme
To investigate the dimension of the degenerate subspace, we start from the degeneracy of the dynamics of the perfect driven Dicke model represented by the toal spin operators in Eq. (37). The angular momentum coupling basis is very useful when it comes to the dynamics described by total spin operators. For the dependence of the degeneracy on the transmon number, we use the eigenstates of the operators $\{\hat{S}_{1}^{2},\hat{S}_{12}^{2},\hat{S}_{13}^{2},...,\hat{S}_{1N}^{2},\hat{S}_{z}\}$, where $\hat{S}_{1n}^{2}=(\hat{S}_{1}+\hat{S}_{2}+...+\hat{S}_{n})^{2}$. Given $S_{n}=\frac{1}{2}$ and $\hat{S}_{1N}=\hat{S}$, the eigenstates can be written as:
@@
If $a$ or $b$ can take $D_{S}$ values, or the degeneracy for the state |$\ket{S,m}$ is $D_{S}$, there should be $D_{S}^{2}$ possible steady-state density operators associated with the total spin $S$. By [1, 63, 94], the expression of $D_{S}$ for $N$ transmons with total spin $S$ reads:
@@
So the total degeneracy (or the possible number of steady states) $N_{ss}$ is:
$N_{ss}=\sum_{S}D_{S}^{2}=\frac{(2N)!}{N!(N+1)!}$ (36)
Due to the factorial in Eq. (35), the degeneracy increases rapidly with $N$. Using Eq. (36) and (35), we calculate the degeneracy and compare it to the results from the diagonalization of the Liouvillian in Fig. 11. The numerical diagonalization of the Liouvillian shows perfect agreement to the theoretical prediction using the angular momentum basis and both grow somewhat faster than exponentially. This justifies our choice of the angular momentum coupling basis to represent the degenerate subspace. In the following discussion, we will use the angular momentum basis to describe the effect of perturbed dynamics in the degenerate subspace.
Figure 11: Comparison between the degeneracy from the diagonalization of the Liouvillian and the theoretical prediction by the angular momentum method, Eqs (36) and (35).
@@ Appendix A Degeneracy of the Liouvillian’s eigenvalues, counting the number of possible steady states
Because the equation of motion for the perfect driven Dicke model in Eq. (37) and (40) does not depend on the angular momentum coupling, any choice of $a$ and $b$ will share the same coefficients $\rho^{S}_{M,M^{\prime}}$, which is the origin of the degeneracy of the possible steady-state density operators. Given total spin $S$ and transmon number $N$, the degeneracy of the angular momentum coupling degrees of freedom $D_{S}$ is shown in Eq. (35).